Mochi:

Taylor wants to bake cookies for his end-of-season football party. The recipe he is using makes 12 cookies per batch and costs $4.50 per batch to make. If Taylor wants his 52 guests to get 2 cookies each, and he eats 2 cookies, how much will it cost Taylor to make the cookies if he only makes whole batches?

Constipated:

whats your answer

xxDeppressionxx:

Do you have any answer choices?

Mochi:

Yes, I tried to edit them in but it didn't work. A. $36.00 B. $42.75 C. $52.50 D. $40.50

jhonyy9:

52 guests to get 2 cookies each so this mean 52*2 = 104 he eats 2 so 104+2 = 106 12 cookies per 1 batch 106 cookies per x batch x = 106*1/12 = 106/12 = 8,83 what mean 9 batches $4,50 the cost per 1 batch x will cost 9 batches x = 9 *4,50 = ? hope helped understandably easy @XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 52 guests to get 2 cookies each so this mean 52*2 = 104 he eats 2 so 104+2 = 106 12 cookies per 1 batch 106 cookies per x batch x = 106*1/12 = 106/12 = 8,83 what mean 9 batches $4,50 the cost per 1 batch x will cost 9 batches x = 9 *4,50 = ? hope helped understandably easy @XioGonz \(\color{#0cbb34}{\text{End of Quote}}\) I agree